4. Use the appropriate compound angle formula to determine the exact value of each expression

a) sin ( pi + Pi over 6 )
b) tan ( Pi over 4 + Pi )
c) cos ( -Pi - pi over 4 )
1
Expert's answer
2020-10-29T20:03:06-0400

a) sin("\pi" )cos("\pi"/6) + cos("\pi" )sin("\pi"/6) = 0 + (-1)(1/2) = -1/2

b) tan("\pi"/4 + "\pi") = (tan("\pi" /4) + tan("\pi")) / (1-tan("\pi"/4)tan("\pi")) = (1 + 0)/(1-0) = 1

c) cos (-)("\pi" +"\pi"/4) = cos("\pi"+ "\pi" /4) = cos("\pi" )cos("\pi"/4) - sin("\pi")sin("\pi"/4) = (-1)(1/ "\sqrt{2}" ) + 0 = -1/"\sqrt{2}"